OpenStudy (anonymous):

What is the slope of the line? y=(5/3)x-5 )

OpenStudy (potatoes.ramu):

Slope =5/3!! From the equation y=mx+c

OpenStudy (anonymous):

And I will give a medal

OpenStudy (studygurl14):

y = mx + b where m = slope Thus... in the equation y = (5/3)x - 5 5/3 = m = slope :)